how do i solve, solve equation, for equation with real solutions, support answer graphically.
(3x-1)^2 = 12

Answer by stanbon(75887)

(3x-1)^2 = 12
----
3x-1 = sqrt(12) or 3x-1 = -sqrt(12)
---
3x-1 = 2sqrt(3) or 3x-1 = -2sqrt(3)
---
3x = 1+2sqrt(3) or 3x = 1-2sqrt(3)
=====================================
Cheers,
Stan H.
==========